OpenStudy (anonymous):

arrange these from least to greatest -3/7 -5/6 -3/5 -7/9

OpenStudy (cwrw238):

convert them all to decimal form - use ur calculator then it is easy to see the differences

OpenStudy (cwrw238):

eg -3/7 = -0.4286

OpenStudy (anonymous):

\[-\frac{ 3 }{ 7 }=-0.42;-\frac{ 5 }{ 6 }=-0.83;-\frac{ 3 }{ 5 }=-0.6;-\frac{ 7 }{ 9 }=-0.77\]  The order is from left to right of the arrow